OpenStudy (anonymous):

. Determine an equation for the asymptote of the graph of y = 2^(x+3)+ 4

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

how?

myininaya (myininaya):

what kindof asymptote?

myininaya (myininaya):

y=4 is a horiztonal asymptote as x gets big f(x) does not get close to anything it just gets really big but if x gets negative big f(x) approaches 4

myininaya (myininaya):

asdjung plug some negative big values of x and see what happens to y

myininaya (myininaya):

what happens if you plug in -700 -800 -900

OpenStudy (anonymous):

how do you do that without a calculator?

myininaya (myininaya):

if you have y=2^x what do you think happens to f(x) as y gets closer to negative infinity

myininaya (myininaya):

f(-5)=2^(-5)=1/32 when you have a negative you will get a fraction

myininaya (myininaya):

so if we plug -9 in we would get something bigger on bottom than 32 right?

myininaya (myininaya):

so 1/(a really big number) is close to zero

OpenStudy (anonymous):

aight i get it now ! thnx

myininaya (myininaya):

ok